OpenStudy (anonymous):

Simplify. Please be sure to show all of your work. 3/8-1/2+5/6

OpenStudy (lalaly):

first you need a comon denominator so 3/8 - 1/2 = 3/8 - (1/2)*(4/4)= 3/8 -4/8 = -1/8 now u hav -1/8 +5/6 = (-1/8)(3/3) + (5/6)(4/4)= -3/24 + 20/24 = 17/24